Tea worth 25.50 per pound is to be mixed with tea worth 29.40 per pound. How many pounds of each should be used to obtain 25 pounds of blend worth 27.00 per pound?
Answer by richwmiller(17219)

25*27=675
x+y=25,
25.5*x+29.4*y=675
x=25-y
25.5*(25-y)+29.4*y=675
637.5-25.5y+29.4*y=675
3.9*y=37.5
y=9.61538462
x=25-y
x=15.3846154 at 25.50 and y=9.61538462 at 29.40
rounded
x=15.4 at 25.50 and y=9.6 at 29.40
check
15.3846154+9.61538462=25
25.5*x+29.4*y=675
25.5*15.3846154+29.4*9.61538462=675
392.307692+282.692308=675
675.0=675
ok